OpenStudy (anonymous):

Choose two axioms that allow 6 + (x + 5) to be written x + 11. commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

Commutative (addition) => x + y = y + x Associative (addition) => x + (y + z) = (x + y) + z